OpenStudy (anonymous):

What is the value of x in the solution to the following system of equations? x − y = −3 x + 3y = 5

OpenStudy (anonymous):

help me pleas ill medal and fan

OpenStudy (anonymous):

Dude, just go to WolframAlpha.com...

OpenStudy (anonymous):

y = x + 3

OpenStudy (campbell_st):

triple the 1st equation then add it to the 2nd equation, that will eliminate y does that make sense

OpenStudy (anonymous):

y = x

OpenStudy (anonymous):

so if i do that it would be 3

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

Is my profile picture ugly?

OpenStudy (michele_laino):

please subtract the second equation from the first one

OpenStudy (campbell_st):

triple every term... is like multiplying every term by 3 3x - 3y = -9 + x + 3y = 5 -------------- what add the 2 equations -3y + 3y = 0 so y is eliminated

OpenStudy (anonymous):

oh i get it

OpenStudy (campbell_st):

so what do you get as the next step

jagr2713 (jagr2713):

@OpiGeode what do you learn when u go on WolframAlpha and or other sites like it?

OpenStudy (anonymous):

wait so now we are solving for x right

OpenStudy (anonymous):

Basically you type in a math problem and it tells you how to solve it and gives you answer along with it.

jagr2713 (jagr2713):

you have to pay for it to give you step by step solutions.

OpenStudy (anonymous):

No..

OpenStudy (campbell_st):

no... before you solve.... just add the 2 equations..... y will disappear... because -3y + 3y = 0 so what is 3x + x = ? and also -9 + 5 = ?

jagr2713 (jagr2713):

Yes you do...

OpenStudy (anonymous):

-9 + 5 = 3

OpenStudy (campbell_st):

well not quite, can you -9 + 5 = 4 what about 3x + x =

OpenStudy (campbell_st):

dang.... -9 + 5 = -4

OpenStudy (anonymous):

oh ok i thought it was -8 minus 5 i didnt see it right

OpenStudy (campbell_st):

so what about the 3x + x = ?

OpenStudy (anonymous):

you have too bring down the 3 right

OpenStudy (campbell_st):

no its 3x + 1x = ?x what is the value of ?

OpenStudy (anonymous):

so its 4

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

great so the problem sets up as 3x - 3y = -9 + x + 3y = 5 -------------- 4x = -4 now solve for x

OpenStudy (anonymous):

-1

OpenStudy (campbell_st):

great, that's it x = -1 all you need to do

OpenStudy (anonymous):

thank you so much

OpenStudy (radar):

You do know that you can subtract as well in the "elimination" process. This is a case where you could just subtract the bottom equation from the top equation thus eliminating the "x" with a result of: -4y = -8 solving for y getting y=2